the length of rectangle is increased by 2 cm. if length is increased by 4cm and breadth is decreased by 3cm. the area remains same. find the length & breadth of rectangle
Answer by CharlesG2(834)

the length of rectangle is increased by 2 cm. if length is increased by 4cm and breadth is decreased by 3cm. the area remains same. find the length & breadth of rectangle
L Length in cm
B breadth (another word for width) in cm
L + 4 cm, B - 3 cm
A=L*B=(L+4) * (B-3)
L*B=L*B-3*L+4*B-12
L*B-L*B+3*L-4*B=-12
3*L-4*B=-12
Thinking what you mean by 1st sentence --> The length of rectangle is increased by 2 cm. --> is that L=B+2
So plugging that in...
3*(B+2)-4*B=-12
3*B+6-4*B=-12
6-B=-12
-B=-18
B=18
so L=20 cm and B=18 cm
verifying: 20*18=360 cm^2
(20+4) * (18-3) = 24*15 = 360 cm^2